Use the Refereaces to access important values if seeded for this question In the laboratory you dissolve 16.6 g of iron(III) acetate in a volumetric flask and add water to a total volume of 125 mL What is the molarity of the solution? 569 What is the concentration of the iron(III) cation? What is the concentration of the acetate anion? Submit Anser Retry Entire Group 9more group attempts remaining 8
media%2F48c%2F48c208f8-a22d-4705-8cf1-f9
Answer #1

1)

a)

Molar mass of Fe(CH3COO)3,

MM = 1*MM(Fe) + 6*MM(C) + 9*MM(H) + 6*MM(O)

= 1*55.85 + 6*12.01 + 9*1.008 + 6*16.0

= 232.982 g/mol

mass(Fe(CH3COO)3)= 16.6 g

number of mol of Fe(CH3COO)3,

n = mass of Fe(CH3COO)3/molar mass of Fe(CH3COO)3

=(16.6 g)/(232.982 g/mol)

= 7.125*10^-2 mol

volume , V = 125 mL

= 0.125 L

Molarity,

M = number of mol / volume in L

= 7.125*10^-2/0.125

= 0.570 M

Answer: 0.570 M

b)

[Fe3+] = [Fe(CH3COO)3] = 0.570 M

Answer: 0.570 M

c)

[CH3COO-] = 3*[Fe(CH3COO)3] = 3*0.570 = 1.71 M

Answer: 1.71 M

Only 1 question at a time please